OpenStudy (anonymous):

Which of the following is the correct factored form of the given equation? 4x^2 - 11x + 6 = 0 (2x + 3)(2x + 2) = 0 (2x - 3)(2x - 2) = 0 (4x - 3)(x - 2) = 0

OpenStudy (anonymous):

If you're unsure you should try multiplying them out. All the information needed is in the following equation (x + a)(x + b) = x^2 + 2ab + ab

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

Factor: 4x^2 - 11x + 6. (1) You may use the factoring AC Method (YouTube) Find 2 numbers b1and b2 that satisfy these 2 conditions: Product b1*b2= ac = 24. Sum (b1 + b2) = b = -11. They are: b1 = -3 and b2 = -8. Replace in the equation (1) the term (-11x) by the 2 terms (-3x) and (-8x) then factor it. y = (x -2)(4x - 3)